Light traveling in air is incident on a pane of glass (n = 1.6) at an angle of 25 degrees relative to the normal. What angle relative to the normal will it have once it is inside the glass?
1
Expert's answer
2019-03-19T12:08:34-0400
"n = \frac{\sin(i)}{\sin(r)}""1.6 = \frac{\sin(25)}{\sin(r)}""r=\sin^{-1}(\frac{\sin(25)}{1.6})"

Thus, i = 15.3 degrees.